OpenStudy (mathisfun13):

would the ratio of 1/7+1/49+1/343 be 7?

OpenStudy (zzr0ck3r):

(1/49)/(1/7) = (1/49)*7 = 1/7

OpenStudy (mathisfun13):

i thought so but when i divide that by 1-1/7 i get 1

OpenStudy (campbell_st):

the denominator 1 - 1/7 = 6/7

OpenStudy (mathisfun13):



OpenStudy (zzr0ck3r):

should be \[\frac{\frac{1}{7}}{1-\frac{1}{7}}\]

OpenStudy (zzr0ck3r):

the ratio is not 7

OpenStudy (zzr0ck3r):

\[\frac{first.term}{1-\frac{second.term}{first.term}}\]

OpenStudy (mathisfun13):

than you would get 1/6

OpenStudy (zzr0ck3r):

correct